# EVE: Towards End-to-End Video Subtitle Extraction with Vision-Language Models

### v1

@@ 5.3 Experimental Results
Table 3: Ablation studies on the proposed EVE (metric: SubER). Through experimental results, the separate projector is necessary for InterleavedVT.
5.4 Ablation Studies
@@
In this paper, we propose an innovative adapter called InterleavedVT, which includes a visual compressor (VC) and a textual region compressor (TRC) for compressing visual features. Additionally, we introduce the separate projector (SP) to project the visual tokens and text-related tokens into the input space of large language models. In this section, we conduct ablation studies on these three modules. The experimental results, as shown in Tab. 3, indicate that introducing either the visual compressor or the textual region compressor alone does not outperform previous methods. However, when both of them are integrated into InterleavedVT, there is a significant improvement in model performance. Compared to only using the textual region compressor, our method achieves a 5.03% and 5.28% improvement in SubER metrics on the positive samples of short video and movie datasets, respectively. Furthermore, introducing separate projectors can further improve the performance. A plausible explanation is that the two types of information focus on different aspects of visual effects, and thus, separate projectors facilitate better alignment of features with LLMs.
6 Limitations and Future Work
Although the proposed EVE have compressed each frame into 26 tokens and supported to input 120 frames for video subtitle extraction, it still have difficulties in solving long videos (e.g., longer than one hour). In future work, we will explore to more effectively compress text-related tokens in video frames to increase the number of image inputs to the model. Furthermore, the proposed EVE can only predict timestamps and texts of subtitles, ignoring the positional information of texts (i.e., bounding box). To support more downstream applications, we will attempt to predict the accurate coordinates of subtitles in each frame.
7 Conclusion
In this paper, we propose an End-to-end Video Subtitle Extraction model, called EVE, which consists of a vision encoder, an adapater module and a large language model. In EVE, a novel adapter, called InterleavedVT, is introduced for effectively compressing visual tokens. To benchmark the video subtitle extraction task, we propose a large bilingual (i.e., Chinese and English) dataset ViSa, including 2.35M videos for pre-training, 230K videos for fine-tuning and 1.6K videos for test. Through experimental results, we observe that the proposed EVE achieves the best performance in both NED and SubER compared with previous methods for video subtitle extraction.
Supplementary Material
8 SRT Format
An SRT file is one of the most popular subtitle file formats for video content. These plain text files include the text of the subtitles in sequence, along with the start and end timestamps. However, an SRT file does not contain any video or audio content. That’s why the file sizes are so small. An SRT file needs to specify the line number, subtitle content, when the text should appear on the video, and when it should disappear to make room for the next line. An example is shown in the following.
Figure 6: An example of SRT files.
9 Hyperparameter Selection
@@
There are two hyperparameters in EVE: the number of queries in the textual region compressor and the number of visual tokens. We have conducted corresponding experiments with different numbers of tokens. When the number of visual tokens (denoted as $N$) and the number of queries in the textual region compressor (denoted as $K$) are set to larger values of ($N$=36 and $K$=20), the performance of EVE is only improved by 0.3% in the NED metric, but the computational cost will increase fourfold. When $N$ and $K$ are set to smaller values ($N$=4 and $K$=1), there is a significant decrease in the performance of EVE, with the NED metric decreasing by nearly 28.6%. In order to balance computational cost and model performance, we ultimately set $N$ and $K$ to 16 and 10, respectively.
10 Comparison with Token Merging Methods
Except the open-sourced subtitle extraction tools and LVLMs, we also compare our method with two typical token Merging methods (T-Selector [30] and ToMe [5]). The results are shown in Tab. 4. We use Qwen2-7B to implement T-Selector and ToMe for subtitle extraction. The experimental results show that although our method EVE-0.5B only adopts a 0.5B language model, it can achieve comparable results to T-Selector and ToMe. When we use Qwen2-7B for EVE, our method can achieve significant performance improvements. For example, compared to ToMe , EVE can achieve a 16.8% performance improvement (in NED) on the Short Videos test dataset. The results demonstrate that the proposed token merging method InterleavedVT can more efficiently extract the subtitle-related information in videos.
Method Short Video Movie
Positive Negative Positive Negative
NED SubER SubER NED SubER SubER
T-Selector [30] 0.801 20.32 23.84 0.731 8.72 19.30
ToMe [5] 0.714 28.35 25.62 0.705 10.79 15.59
EVE-0.5B 0.759 22.82 18.36 0.685 5.10 17.06
EVE-7B 0.882 15.16 13.92 0.808 1.64 10.15
Table 4: Performance comparison with two typical token merging methods.
11 Visualization for Textual Region Compressor
In the proposed EVE, the textual region compressor is introduced to perceive text-related information in each frame. To verify the effectiveness of this module, we have visualized the attention map of the textual region compressor in Fig. 7. In the visualization, we observe that this module can indeed focus on the text region in each frame, which demonstrates the effectiveness of the proposed textual region compressor.
Figure 7: The visualization for the textual region compressor.
12 More Statistical Information of ViSa
The detail information of SFT dataset have shown in the main text. In this section, we show some statistical information of the pre-training data in Fig. 8. Most of pre-training videos contain less than 10 subtitles and their duration is between 11-20 seconds.
13 More Result Visualizations
We additionally visualize two comparison examples. Fig.9 and Fig.10 show the short video and movie examples, respectively. Through these two examples, we observe that Qwen2-VL and MiniCPM tend to describe the video content rather than predicting subtitles in SRT format. In addition, the cloud services can hardly extract bilingual subtitles, especially in movie videos. The ASR-based method PyVideoTrans also cannot achieve ideal performance since it only takes audio as input, which is not fully aligned with visual information. Compared with previous methods, the proposed EVE can better predict texts and timestamps of subtitles.
Figure 8: The statistics information of ViSa’s pre-training dataset.
Figure 9: A comparison example of short video.
Figure 10: A comparison example of movie.

### v2

@@ V-C1 Evaluating Existing Methods
V-C2 Comparison with Existing Methods
In this paper, we establish two models with different LLMs. To achieve better performance in video subtitle extraction, we utilize Qwen2-7B to construct EVE-7B. Through the experimental results in Table III, EVE-7B achieves the best performance compared with three types of video subtitle extraction methods. For example, on the short video dataset, EVE-7B outperforms OpenVision by 11.1% in the metric of NED. Considering that the EVE-7B will cost more computing resources in practical applications, we establish a light-weighted model EVE-0.5B with Qwen2-0.5B. The experimental results in Table III demonstrate that despite using the same dataset for training, EVE-0.5B still cannot achieve the same performance as EVE-7B. However, benefiting from the proposed framework and training strategy, it can still surpass most of existing methods.
Figure 6: An example of subtitle extraction for different methods. Gray and Yellow indicate the missing and repetition prediction problem, respectively. Red represents the format of the answer is incorrect. Green represent the correct prediction content.
Figure 7: The visualization for the temporal subtitle token querying using heatmap.
Figure 8: Visualization of subtitle extraction results with timestamps from a movie clip using MiniCPM, GhostCut, VideOCR, Qwen2-VL, and the proposed EVE.
V-D Comparison with Token Compression Methods
Beyond comparisons with existing subtitle extraction tools and LVLM-based baselines, we further evaluate the proposed S3 module against two representative token compression approaches, T-Selector [wang2025elysium] and ToMe [bolyatoken], as summarized in Table IV. As mentioned in the Related Work section, these methods are functionally similar to the proposed S3 module, as they all aim to extract key information to feed into the LLM. Therefore, we conduct this experiment for comparison. Both methods are re-implemented using the Qwen2-7B backbone to ensure a fair comparison under identical conditions. The results reveal that despite utilizing a smaller 0.5B language model, EVE-0.5B achieves performance on par with or even surpasses these token compression baselines, validating the efficiency of our design. When scaled to EVE-7B, the performance improves substantially across all benchmarks. For instance, compared to ToMe, EVE achieves a 16.8% improvement in NED on the Short Videos test set, demonstrating stronger temporal alignment and finer-grained token utilization. This consistent advantage highlights the effectiveness of our proposed S3 module, which interleaves vision and text representations to selectively preserve semantically relevant tokens while filtering out redundant ones. Unlike conventional compression schemes that rely purely on token similarity, S3 module adaptively integrates multimodal cues, enabling more precise identification of subtitle-relevant regions and enhancing overall extraction robustness across both short and long-form videos.
V-E Ablation Studies
@@
In this paper, we propose an innovative adapter called S3 module, which includes a Spatial Semantic Context Aggregate (SSCA) and a Temporal Subtitle Token Query (TSTQ). Additionally, we introduce the separate projector (SP) to project the global visual tokens and text-related tokens into the input space of large language models. In this section, we conduct ablation studies on these three modules. The experimental results, as shown in Table V, indicate that introducing either the SSCA or the TSTQ alone does not outperform previous methods. However, when both of them are integrated into S3 module, there is a significant improvement in model performance. Compared to only using the TSTQ, our method achieves a 5.03% and 5.28% improvement in SubER metrics on the positive samples of short video and movie datasets, respectively. Furthermore, introducing separate projectors can further improve the performance. A plausible explanation is that the two types of information focus on different aspects of visual effects, and thus, separate projectors facilitate better alignment of features with LLMs.
V-F Hyperparameter Selection
@@
There are two hyperparameters in EVE: the number of queries in the TSTQ and the number of visual tokens after SSCA. We have conducted corresponding experiments with different numbers of tokens. When the number of visual tokens (denoted as $N$) and the number of queries in the TSTQ (denoted as $K$) are set to larger values of (NN=36 and KK=20), the performance of EVE is only improved by 0.3% in the NED metric, but the computational cost will increase significantly, due to more tokens being reserved for the LLM. When $N$ and $K$ are set to smaller values (NN=4 and KK=1), there is a significant decrease in the performance of EVE, with the NED metric decreasing by nearly 28.6%. In order to balance computational cost and model performance, we ultimately set $N$ and $K$ to 16 and 10, respectively.
V-G Visualization for Temporal Subtitle Token Query
To better understand the role of the Temporal Subtitle Token Query in EVE, we visualize its attention responses across representative frames, as shown in Figure 7. The upper row depicts the original video frames, while the lower row illustrates the corresponding attention heatmaps. We observe that the TSTQ consistently activates in regions containing textual elements, such as interface buttons, titles, and overlaid text cues. This selective activation indicates that the module effectively distinguishes text-related areas from irrelevant visual content, enabling more efficient information compression and representation. Moreover, as the video progresses, the attention maps dynamically adjust to newly appearing text, demonstrating strong adaptability to temporal variations. These visualizations validate that the TSTQ successfully enhances the model’s perception of text-bearing regions and contributes to more precise multimodal understanding within EVE.
V-H Visualization for Video Subtitle Extraction
In Figure 6, we compare PyVideoTrans [pyvideotrans], OpenVision [openvision], Qwen2-VL [wang2024qwen2], and the proposed EVE for subtitle extraction. PyVideoTrans omits critical dialogue segments, while OpenVision introduces redundancy through repeated outputs. Qwen2-VL diverges from the task objective, producing scene-level descriptions rather than precise transcripts. By contrast, EVE generates temporally aligned, semantically faithful, and duplication-free subtitles. For example, the utterance “Oh, have they gone off again, have they?” is recovered accurately and uniquely. These results confirm EVE’s robustness and highlight its effectiveness as a reliable framework for multimodal subtitle generation.
In Figure 8, we further compare MiniCPM  [yao2024minicpm], GhostCut [ghost], VideOCR [videocr], Qwen2-VL [wang2024qwen2], and EVE for video subtitle extraction. MiniCPM fails to perform transcription and instead outputs verbose scene-level descriptions. GhostCut captures most subtitle lines but exhibits minor temporal misalignment and redundant repetitions. VideOCR shows high recognition quality but suffers from token concatenation issues such as “It’s ok, Truman. I understand.” caused by inaccurate text segmentation. Qwen2-VL generates only a single incomplete line, failing to cover the entire dialogue sequence. In contrast, EVE delivers temporally consistent, structurally complete, and duplication-free subtitles. For instance, the utterance “It’s ok, Truman. I understand.” is precisely aligned and clearly formatted, demonstrating EVE’s strong capability in maintaining accurate temporal correspondence and structural completeness. Moreover, the visualization results indicate that the timestamps predicted by EVE are closer to the ground truth compared to other methods.
VI Conclusion
In this paper, we propose an End-to-end Video Subtitle Extraction model, called EVE, which consists of a vision encoder, an adapter module and a large language model. In EVE, a novel adapter called the S3 module has been introduced, which is able to model salient subtitle representations while retaining only a minimal number of tokens. To benchmark the video subtitle extraction task, we propose a large bilingual (i.e., Chinese and English) dataset ViSa, including 2.35M videos for pre-training, 230K videos for fine-tuning and 1.6K videos for test. Through experimental results, we observe that the proposed EVE achieves the best performance in both NED and SubER compared with previous methods for video subtitle extraction. Despite its strong performance, EVE still faces two limitations. First, although each frame is compressed into only 26 tokens, the maximum input length remains insufficient for very long videos, such as those exceeding one hour. Second, EVE predicts only subtitle texts and timestamps, without providing spatial coordinates of subtitle regions. Future work will explore more efficient token compression strategies to support longer input sequences and extend EVE to localize subtitle positions for broader downstream applications.